OpenStudy (anonymous):

if I assume that cos(x)+5sin(x)=23/5 and 2cos(x)+sin(x)=10/5 how would I find the exact numeric value of sec(x) and tan(x)?

OpenStudy (tkhunny):

Can you solve this system? x + 5y = 23 2x + y = 10

OpenStudy (jdoe0001):

yeah is a system of equations

OpenStudy (tkhunny):

Excellent. Apply the same principles. cos(x)+5sin(x)=23/5 and 2cos(x)+sin(x)=10/5 From the first, cos(x) = 23/5 - 5sin(x) Substitute into the second. 2[23/5 - 5sin(x)]+sin(x)=10/5 This leads to an exact value for sin(x) and you're almost done.